OpenStudy (anonymous):

Please Help! Will give a medal! :) How many REAL number solutions does the equation 7x2 - 5x + 1 = 0 have?

OpenStudy (whpalmer4):

\[7x^2-5x+1=0\]is a quadratic of the form \[ax^2+bx+c=0\]The discriminant \(\Delta = b^2-4ac\) will characterize the solutions: If \(\Delta = 0\), a perfect square with 1 real solution if \(\Delta > 0\), two real roots if \(\Delta < 0\), two complex roots in a conjugate pair \((a\pm bi)\)